Physical chemistry question
For the reaction: creatine phosphate + H2O right
Answer #1

Apply:

Q = [products ] /[reactas]

Q = (7.8*10^-1)(9.2*10^-1)/(1.2*10^-7) = 5980000 = 5.98 *10^7

Q for std. state...

dG = -RT*lnK

K = exp(37600/8.314/310) = 2166632 = 2.16*10^7

since Q > K, this favours forward direction